# Quantum Blackwell’s Ordering and Differential Privacy

### v1

@@ Abstract
We develop a framework for quantum differential privacy (QDP) based on quantum hypothesis testing and Blackwell’s ordering. This approach characterizes $(\varepsilon,\delta)$-QDP via hypothesis testing divergences and identifies the most informative quantum state pairs under privacy constraints. We apply this to analyze the stability of quantum learning algorithms, generalizing classical results to the case $\delta>0$. Additionally, we study privatized quantum parameter estimation, deriving tight bounds on the quantum Fisher information under QDP. Finally, we establish near-optimal contraction bounds for differentially private quantum channels with respect to the hockey-stick divergence.
@@ I Introduction
A fundamental challenge in modern machine learning is the trade-off between privacy and information extraction. In this work, we explicitly treat both sides: privacy (ensuring that algorithmic outputs do not reveal significant information about the input data of the respondents) and the investigator’s goal to extract as much useful information as possible from data for accurate learning and estimation. With the rapid advancement of machine learning, a key concern is about ensuring the privacy of learning algorithms, meaning that their outputs should not reveal significant information about the input data. Differential privacy (DP) provides a rigorous mathematical framework to balance these opposing requirements. Accordingly, we structure our contributions in three steps: first step (privacy), second step (information extraction under privacy constraints), and third step, the quantum channel setup, where the situation is more complicated, and we mark the transition to each step explicitly in the text.
First step: privacy. This step develops the privacy side of the trade-off from the respondent’s perspective by studying the stability [1, 2] of learning algorithms. From the respondent’s viewpoint, privacy means that the inclusion or exclusion of their individual data should not materially affect the mechanism’s output, so that they can contribute data without fear of singled-out inference. An algorithm is considered stable if its output does not change drastically when a single respondent’s data is changed; this point-wise insensitivity is precisely the respondent-centric guarantee we seek. Differential privacy $($$\varepsilon,\delta$)-DP) [3, 4] formalizes this requirement as a strong, mathematically precise form of stability: by ensuring that the output states (distributions) corresponding to any two neighboring datasets are close, an $(\epsilon,\delta)$-DP algorithm guarantees that no respondent can be reliably distinguished from the output alone [5].
In the classical case, the stability of learning algorithms was studied in [6] only for the pure $\varepsilon$-DP ($\delta=0$) case. The main aim of this manuscript is to generalize the result of [6] to the quantum learning scenario for $(\varepsilon,\delta)$-DP. To achieve this, we build a framework for studying quantum differentially private mechanisms $($$\varepsilon,\delta$)-DP). This framework is inspired by [7] and is related to the problem of hypothesis testing (see for example, [8]). The main aim of differential privacy is to ensure that the outputs of a private mechanism corresponding to two neighboring inputs are hard to distinguish. To understand this, consider the following scenario. Suppose inside a database of datasets, we have two datasets $S$ and $T$ such that there is exactly one individual, say $A$ is present in set $S$ and not in $T$, i.e., $S\setminus(S\cap T)=\{A\}$ and suppose there is an oracle (randomized) which is attached to the database. From the perspective of an attacker, the only way it can break the privacy of this oracle is by performing a hypothesis test on the output of the oracle to detect whether the individual $A$ is a member of $S$ or not. From this discussion, it implies that the more informative the oracle is about the datasets, the less private it is.
A notion for comparing the “informativeness” between two mechanisms (oracles) was first studied by Blackwell [9]. In [10, Theorem 12.2.2], he showed that if a mechanism (oracle) is more informative as compared to the other mechanism, then there exists a Markov kernel (channel) which maps the more informative one to the less informative one. A quantum version of [10, Theorem 12.3.1] was proved by Buscemi in [11], wherein he defined a notion of ordering between quantum channels based on a certain definition of informativeness. Using this ordering, Buscemi showed that there exists a statistical morphism from the more informative quantum channel to the lesser one.
[…]
Using Lemma 1 (a quantum version of [10, Theorem 12.4.2]), we show that there exists a worst case $(\varepsilon,\delta)$-DP pair of quantum states which can be mapped to any other $(\varepsilon,\delta)$-DP pair of states by applying a CP-TP map. Thus, this version of the quantum Blackwell theorem (Lemma 1), along with this worst case $(\varepsilon,\delta)$-DP pair, provides us with a powerful machinery to analyze the stability of quantum differentially private $($$\varepsilon,\delta$)-DP) learning algorithms.
Second step: information extraction under privacy constraints. We now analyze, under the $(\varepsilon,\delta)$-DP constraints, how much useful information an investigator can still extract. That is, we investigate the statistical inference of quantum privatized parameter using our quantum version of Blackwell’s theorem (Lemma 1). This is a common task where we want to guess a hidden number or parameter from data that has been made private. How well the investigator can guess this number is limited by how much information the privacy method keeps. This limit is measured by a value called the Fisher information. In the quantum case, we use a related measure called the Symmetric Logarithmic Derivative (SLD) Fisher information to judge performance [12]. By using the idea of our “weakest” (most informative) private method, we can figure out the exact maximum SLD Fisher information the investigator can get while still respecting the $(\varepsilon,\delta)$-DP rule. This work improves on past studies that only looked at the simpler case where $\delta=0$ [13], and sets a hard limit on how accurate quantum estimations can be under the privacy condition.
Third step: We have two kinds of extensions. In the privacy of the quantum learning framework, we can relax the $(\varepsilon,\delta)$-DP condition by trusting the party, the data processor, who performs the learning algorithm. That is, motivated by the quantum learning framework of [14, 15], we propose a framework for a quantum learning algorithm which is “$1$-neighbor" $(\varepsilon,\delta)$-DP. For this proposed learning framework, we study its stability by obtaining an upper-bound on the Holevo information (between the training data and the quantum output of the algorithm). Further, we show that for $\delta=0$, our upper-bound recovers [6, Proposition 2] in the classical case.
@@
Another interesting problem is to analyze quantum channels which are $(\varepsilon,\delta)$-DP. Its classical version has been extensively studied in the literature via the contraction of output divergence-measured with respect to the output distributions induced by $(\varepsilon,\delta)$-DP channels-in comparison to the input divergence, which is defined over the input distributions of these channels; see, for example, [16], [17], and the references therein. This contraction is studied using an integral representation which involves the hockey stick divergence [18]. We leverage our framework to first establish an almost tight upper-bound on the contraction coefficient of any $(\varepsilon,\delta)$-DP CPTP map with respect to the hockey stick divergence.
In the process of studying the contraction for divergences, we observe that in the classical case, there exists a pure $(\varepsilon+\log(1/1-\delta))$-DP pair of distributions which can be obtained by truncating the original $(\varepsilon,\delta)$-DP pair. Further, we show that this truncated pair is not far ($O(\delta)$, in $L_{1}$ distance) from the original pair. We use this truncated pair to obtain a meaningful upper-bound on the relative entropy (when it is well-defined) of the original $(\varepsilon,\delta)$-DP pair of distributions using the continuity of the relative entropy. This upper-bound recovers the known result for $\delta=0,$ [19, 7] as a special case.
Moreover, these contraction-based analysis serve to connect the channel-level action with both privacy guarantees and investigator-agnostic limits on information extraction, by bounding how much input distinctions can be suppressed at the output, we obtain direct, worst-case control of output distinguishability (useful for hypothesis-testing-based privacy statements) and of output information quantities (useful for bounding mutual information, Fisher-information-based estimation limits).
@@
Blackwell’s Informativeness
Theorem
[10, Theorem 12.4.1 and 12.4.2]
(Proposition 1 in this manuscript.)
Lemma 1 in this manuscript.
Existence of Weakest
DP Mechanism
[19, Theorem 5]
[19, Theorem 18]
[13, Theorem 2]
Lemma 4
of this manuscript.
Data Processing based
upper-bounds using Weakest
DP Mechanism
[7, Lemma D.8]
Corollary 3
of this manuscript.
Corollary 2
of this manuscript.
Privatized Parameter
Estimation
[19, Theorem 18]
[19, Theorem 18]
[13, Section IIA]
Theorem 3
of this manuscript.
Stability Upper-bounds
of Private Learning
Algorithms
[6, Proposition 2]
Proposition 2
of this manuscript.
Equations (i)),
(ii)) and (iii)) of
this manuscript.
Theorem 5
of this manuscript.
Theorem 5
of this manuscript.
Privatized Contraction
Coefficient of Trace Distance
[19, Corollary 11]
[20, Theorem 22]
[20, Theorem 55]
Contraction Coefficient
of Hockey-stick Divergence
[17, Theorem 1]
Corollary 9
of this manuscript.
[20, Theorem 1]
Lemma 7
of this manuscript.
Contraction Coefficient based
Upper-bounds on Relative entropy
of LDP channels
Theorem 6
of this manuscript.
[20, Proposition 3]
@@ I-A Organization of this Manuscript and our Contributions
  • In Section III, we develop a quantum generalization of Blackwell’s theorem, which establishes an ordering of informativeness for pairs of quantum states. In Lemma 1, we show that one pair is less informative than the other if and only if there exists a completely positive map connecting them. This forms the foundation for comparing privacy mechanisms in the quantum setting.
  • In Section V, we prove Lemma 3, which shows that a pair of quantum states is $(\varepsilon,\delta)$-differentially private if and only if $D^{\alpha}_{H}{(\rho\|\sigma)}\leq\max\left\{1-\delta-e^{\varepsilon}\alpha,e^{-\varepsilon}(1-\alpha),0\right\}$ for any $\alpha\in[0,1],$ where $D^{\alpha}_{H}{(\rho\|\sigma)}$ is the hypothesis testing divergence between $\rho$ and $\sigma.$ Using this in Lemma 4, we identify an explicit “weakest” (most-informative) pair of $(\varepsilon,\delta)$-DP quantum states that dominates all other.
  • In Section VI, we apply our framework to quantum parameter estimation under privacy constraints. In particular, in Theorem 3, we derive the exact maximum SLD Fisher information achievable by any $(\varepsilon,\delta)$-DP mechanism.
  • In Section VII, we study the stability of $(\varepsilon,\delta)$-DP learning algorithms. We first discuss a framework for differentially private learning algorithms. Then, using this framework, in Theorem 5, we study their stability by deriving upper-bound on the Holevo information.
  • In Section VIII, we analyze the contraction of divergences under differentially private channels. In Lemma 7, we obtain nearly tight upper and lower bounds on the contraction coefficient of $(\varepsilon,\delta)$-LDP channels with respect to the quantum hockey-stick divergence. We then use this to derive a novel upper-bound on the relative entropy for classical $(\varepsilon,\delta)$-LDP channels in Theorem 6 via its integral representation [18].

### v3

# Quantum Information Ordering and Differential Privacy

@@ Abstract
We study quantum differential privacy (QDP) by defining a notion of the order of informativeness between pairs of quantum states. In particular, we show that if the hypothesis testing divergence of one pair dominates over that of the other pair, then this dominance holds for every $f$-divergence. This approach completely characterizes $(\varepsilon,\delta)$-QDP mechanisms by identifying the most informative $(\varepsilon,\delta)$-DP quantum state pairs. We apply this to study precise limits for privatized hypothesis testing and privatized quantum parameter estimation, including tight upper-bounds on the quantum Fisher information under QDP. Finally, we establish near-optimal contraction bounds for differentially private quantum channels with respect to the hockey-stick divergence.
@@ I Introduction
The trade-off between extracting useful insights by an investigator and preserving a respondent’s privacy lies at the heart of modern quantum machine learning. While classical Differential Privacy (DP) [13, 14] provides a rigorous framework to limit adversarial inference via hypothesis testing, extending these guarantees to the quantum regime presents unique challenges. Recent advances in Quantum Differential Privacy (QDP) [36] have made non-trivial attempts [12, 18, 1, 25, 26] to understand differentially private mechanisms in the quantum setting. In this manuscript, we advance the understanding of quantum differential privacy by grounding QDP in the geometry of asymmetric hypothesis testing. We establish a universal information ordering for differentially private mechanisms, proving that the entire privacy regime can be analyzed by a single “most-informative” pair of quantum states. Leveraging this, we derive exact, mechanism-agnostic limits for different statistical tasks: the optimal error exponent in hypothesis testing, the maximum achievable Fisher information for parameter estimation, and the fundamental contraction of distinguishability under private quantum channels.
Framework for Differential Privacy.— In the quantum setting, differential privacy constrains the output which the quantum algorithms produce. Formally, consider a scenario where a respondent prepares a quantum system in state $\tau\in\{\rho,\sigma\}$. The pair $(\rho,\sigma)$ satisfies $(\varepsilon,\delta)$-DP if, for any general quantum measurement (POVM) performed by an investigator, the probability distributions of the measurement outcomes are statistically close. Thus, as defined in [36] for fixed $\varepsilon\geq 0$ and $\delta\in[0,1]$, $(\rho,\sigma)$ is $(\varepsilon,\delta)$-DP if they satisfy the following.
Definition 1.
For every POVM measurement $0\preceq\Lambda\preceq\mathbb{I}$, the following holds:
For the case when $\delta=0$, any value of $\varepsilon$ satisfying Eq. (1) must be at least $\max\{D_{\max}(\rho\|\sigma),\,D_{\max}(\sigma\|\rho)\},$ where $D_{\max}$ denotes the max-relative entropy. As a consequence, the required $\varepsilon$ is typically very large, which severely weakens privacy as discussed in the section below. To mitigate this issue, one introduces a positive additive constant $\delta$, allowing the permissible value of $\varepsilon$ to be significantly reduced. Further in Eq. (1), if we consider only binary POVMs then it motivates us to analyze QDP from the point of view of binary quantum hypothesis testing.
The Operational Scenario: The Respondent and the Investigator.— This manuscript is based on the strategic interaction between two distinct parties: the Respondent (the data owner) and the Investigator (the adversary).
The Respondent serves as the gatekeeper of a sensitive database containing potential datasets $S$ and $T$, which differ solely by the presence of a single individual $A$, such that $S\setminus(S\cap T)=\{A\}$. To make data accessible while maintaining confidentiality, the Respondent instantiates a statistical oracle (the mechanism). This oracle operates as a conditional distribution $\mathbb{P}^{(D)}(O\mid I)$—or in the quantum setting, a quantum channel—generating an output $O$ for a specific query $I$ based on the underlying dataset $D\in\{S,T\}$. The Respondent’s primary objective is to configure this oracle to satisfy strict differential privacy constraints, effectively masking the specific identity of the dataset.
The Investigator operates as the Adversary in this setting. They interact with the Respondent’s oracle by issuing queries and processing the resulting outputs. Their goal is to execute a binary hypothesis test to breach privacy: specifically, to distinguish whether the underlying dataset is $S$ or $T$, thereby revealing the presence of individual $A$. The central thesis of this work is that the “informativeness” of the Respondent’s oracle dictates the success of the Investigator: the more useful the oracle is for legitimate statistical utility, the more vulnerable it is to this adversarial attack. We structure our contributions in three steps corresponding to this dynamic: the privacy limit (comparing the Respondent’s oracles), the adversarial extraction limit (the Investigator’s power), and the channel analysis (the properties of the Respondent’s defense mechanism).
First step: Privacy. The primary objective of differential privacy is to ensure that the outputs of the oracle corresponding to two neighboring inputs are difficult for the Investigator to distinguish. This implies a direct inverse relationship: the more informative the oracle is about the datasets, the less private it is. To formalize this, we require a rigorous method to compare the “informativeness” of different oracles. We consider a pair of oracles modeled as $\mathbb{P}^{(D)}(O\mid I)$ and $\mathbb{Q}^{(D)}(O\mid I)$. We say that $\mathbb{P}^{(D)}(O\mid I)$ is more informative than $\mathbb{Q}^{(D)}(O\mid I)$ (represented as the order $\mathbb{P}^{(D)}(O\mid I)\succeq\mathbb{Q}^{(D)}(O\mid I)$ or simply $\mathbb{P}^{(D)}(O\mid I)$ dominates $\mathbb{Q}^{(D)}(O\mid I)$) if the pair $\{\mathbb{P}^{(S)}(O\mid I),\mathbb{P}^{(T)}(O\mid I)\}$ is statistically easier to distinguish as compared to the pair $\{\mathbb{Q}^{(S)}(O\mid I),\mathbb{Q}^{(T)}(O\mid I)\}$.
In $1951$, Blackwell [4, 3] formalized this notion of comparing “informativeness” for classical statistical experiments. He showed that $\mathbb{P}^{(D)}(O\mid I)\succeq\mathbb{Q}^{(D)}(O\mid I)$ if and only if there exists a Markov kernel (stochastic map) $\mathcal{M}$ such that $\mathbb{Q}^{(S)}(O\mid I)=\mathcal{M}(\mathbb{P}^{(S)}(O\mid I))$ and $\mathbb{Q}^{(T)}(O\mid I)=\mathcal{M}(\mathbb{P}^{(T)}(O\mid I))$. Further, using the data-processing inequality (monotonicity property) we have $D_{f}(\mathbb{P}^{(S)}(O\mid I)\|\mathbb{P}^{(T)}(O\mid I))\geq D_{f}(\mathbb{Q}^{(S)}(O\mid I)\|\mathbb{Q}^{(T)}(O\mid I))$, where $D_{f}$ is the $f$-divergence [29]. For the case when these oracles are differentially private (with the same privacy parameters), it is obvious that the above monotonicity would help us obtain security bounds of the less informative in terms of the one that is more informative.
Likewise, in the quantum domain, we consider two quantum oracles whose outputs are $\{\rho^{S},\rho^{T}\}$ and $\{\sigma^{S},\sigma^{T}\}$. We say that $\{\rho^{S},\rho^{T}\}$ is more informative than $\{\sigma^{S},\sigma^{T}\}$ if the former is statistically easier for the Investigator to distinguish. However, a significant theoretical hurdle arises here: unlike the classical case, it follows from [15] and [23] that the statistical ordering $\{\rho^{S},\rho^{T}\}\succeq\{\sigma^{S},\sigma^{T}\}$ does not imply the existence of a completely positive trace-preserving (CP-TP) map—or even a positive trace-preserving map—that transforms the former into the latter. As a consequence, we cannot rely on the standard data-processing inequalities of quantum $f$-divergences [19] to transfer bounds. We overcome this in Theorem 2 by proving that an explicit channel (a CP-TP map) construction is unnecessary. We show that if the quantum hockey stick divergence [31] between the oracle outputs $\{\rho^{S},\rho^{T}\}$ dominates $\{\sigma^{S},\sigma^{T}\}$ for every parameter, then the inequality $D_{f}(\rho^{S}\|\rho^{T})\geq D_{f}(\sigma^{S}\|\sigma^{T})$ holds universally. This result allows us to define a “worst-case” (most informative) pair of states that characterizes the upper bound on the Investigator’s adversarial capability.
Second step: Information extraction under privacy constraints. We now analyze, under the $(\varepsilon,\delta)$-DP constraints, the fundamental limits of the Investigator’s ability to extract information from Respondent’s data. Here, we treat the Investigator as an adversary whose goal is to infer the hidden parameter or state. We investigate this statistical inference using our quantum $f$-divergence dominance result (Theorem 2). The Investigator’s success in guessing the hidden parameter is strictly limited by the privacy constraints imposed by the mechanism. We examine this limit through two adversarial viewpoints: binary hypothesis testing and parameter estimation.
In the former scenario, we begin with the one-shot setting, where the Investigator attempts to distinguish between two nearby sources $\rho_{\theta_{0}}$ and $\rho_{\theta_{1}}$ (parametrized by two private parameters $\theta_{1}$ and $\theta_{2}$) after their encoding has been privatized through an $(\varepsilon,\delta)$-DP mechanism. This scenario exposes the operational limit of privacy breaches. We show that, among all $(\varepsilon,\delta)$-DP mechanisms, the Investigator’s optimal performance—quantified by the hypothesis testing divergence—is maximized by the “weakest” (most informative) privatized mechanism permitted by the definition. This optimality carries over to KL, hockey stick Rényi (see Definition 12), and measured Rényi divergences. We then extend the analysis to the asymptotic regime and establish that this same privatized mechanism determines the optimal error exponent. This yields a unified characterization: the privacy requirement compresses the geometry of the state space, imposing a fundamental ceiling on the Investigator’s probability of successfully distinguishing hypotheses.
In the regime of parameter estimation, the limit of the Investigator’s ability to pinpoint the private parameter is measured by the Fisher information. In the quantum case, we use the Symmetric Logarithmic Derivative (SLD) Fisher information [17]. By leveraging the idea of our “weakest” (most informative) private method, we determine the exact maximum SLD Fisher information the Investigator can obtain while the mechanism respects the $(\varepsilon,\delta)$-DP rule. This result improves on past studies that only looked at the simpler case where $\delta=0$ [34], and sets a hard physical limit on how accurately an adversarial Investigator can estimate sensitive parameters under the privacy condition.
@@
Third step: The Quantum Channel. Finally, we analyze the mechanism itself by studying quantum channels which are $(\varepsilon,\delta)$-DP. From the adversarial perspective, the channel acts as a defense mechanism that suppresses the distinguishability of inputs. We quantify this suppression via the contraction of divergence—comparing the divergence of the output distributions induced by $(\varepsilon,\delta)$-DP channels to the divergence of the input states; see, for example, [2], [35], and the references therein. This contraction is studied using an integral representation involving the hockey stick divergence [29]. We leverage our framework to first establish an almost tight upper bound on the contraction coefficient of any $(\varepsilon,\delta)$-DP CP-TP map with respect to the hockey stick divergence.
A crucial technical contribution in this step addresses the difficulty of the $\delta>0$ regime. The additive $\delta$ term typically complicates the analysis of entropic quantities. We observe that, similar to the classical case, there exists a pure $(\varepsilon+\log(1/1-\delta))$-DP pair of distributions that can be obtained by truncating the original $(\varepsilon,\delta)$-DP pair. We formally extend this to the quantum domain, showing that this truncated pair is close ($O(\delta)$ in $L_{1}$ distance) to the original pair. We use this truncated pair to obtain a meaningful upper bound on the relative entropy of the original $(\varepsilon,\delta)$-DP pair using the continuity of relative entropy. This upper bound recovers the known result for $\delta=0$ [22, 11] as a special case. Moreover, these contraction-based analyses serve to connect the channel-level action with both privacy guarantees and the limits on adversarial extraction. By bounding how much input distinctions are suppressed at the output, we obtain direct, worst-case control of output distinguishability (useful for bounding the success of hypothesis-testing attacks) and of output information quantities (limiting the precision of Fisher-information-based estimation).
@@
Existence of Weakest DP Mechanism [22, Theorem 5] [22, Theorem 18] [34, Theorem 2] Lemma 3 of this manuscript.
Data Processing based upper-bounds using Weakest DP Mechanism [11, Lemma D.8] Corollary 2 of this manuscript. Theorem 2 of this manuscript.
Privatized Hypothesis Testing [22, Theorem 18] [22, Theorem 18] [34, Section IIB] Theorem 3 and Corollary 4 of this manuscript.
Privatized Parameter Estimation [22, Theorem 18] [22, Theorem 18] [34, Section IIA] Theorem 4 of this manuscript.
Privatized Contraction Coefficient of Trace Distance [22, Corollary 11] [26, Theorem $2$] [26, Theorem $5$]
Contraction Coefficient of Hockey-stick Divergence [35, Theorem 1] Corollary 5 of this manuscript. [26, Theorem 1] Lemma 5 of this manuscript.
Contraction Coefficient based Upper-bounds on Relative entropy of LDP channels Theorem 5 of this manuscript. [26, Proposition 3]
@@ I-A Organization of this Manuscript and our Contributions
  • In Section III, we provide a geometric characterization of the privacy region of $(\varepsilon,\delta)$-QDP and provide a construction of the weakest (most informative) $(\varepsilon,\delta)$-DP pair of quantum states.
  • In Section IV, we first define a notion of order of informativeness between two pairs of quantum states via quantum hypothesis testing divergence. Using this ordering, in Theorem 1, we give an $f$-divergence dominance of a more informative pair of quantum states over a lesser informative pair of quantum states. Further, in Lemma 2, we show that a pair of quantum states is $(\varepsilon,\delta)$-differentially private if and only if $D^{\alpha}_{H}{(\rho\|\sigma)}\leq-\log\max\left\{1-\delta-e^{\varepsilon}\alpha,e^{-\varepsilon}(1-\delta-\alpha),0\right\}$ for any $\alpha\in[0,1],$ where $D^{\alpha}_{H}{(\rho\|\sigma)}$ is the hypothesis testing divergence between $\rho$ and $\sigma.$ Using this in Lemma 3, we identify an explicit “weakest” (most-informative) pair of $(\varepsilon,\delta)$-DP quantum states that dominates all others. Finally, in Theorem 2, we show that the $f$-divergence of any pair of $(\varepsilon,\delta)$-DP quantum states can be upper-bounded by that of the “weakest” (most-informative) pair of $(\varepsilon,\delta)$-DP quantum states.
  • In Section V, we apply our framework to the regime of the quantum hypothesis testing and the quantum parameter estimation under privacy constraints. In particular, in Theorem 3, we show that among all $(\varepsilon,\delta)$-DP mechanisms, the hypothesis testing divergence is maximized by the weakest (most informative) privatized mechanism. Further, in Theorem 4, we derive the exact maximum SLD Fisher information achievable by any $(\varepsilon,\delta)$-DP mechanism.
  • In Section VI, we analyze the contraction of divergences under differentially private channels. In Lemma 5, we obtain nearly tight upper and lower bounds on the contraction coefficient of $(\varepsilon,\delta)$-LDP channels with respect to the quantum hockey-stick divergence. We then use this to derive a novel upper-bound on the relative entropy for classical $(\varepsilon,\delta)$-LDP channels in Theorem 5 via its integral representation [29].